an airplane traveling 280 miles per hour was reported over an airfield. Another plane took off from the airfield 6 minutes later and followed the first plane. The second plane's average speed was 320 miles per hour. Assuming the planes did not change speed and direction, how far from the airfield did the second plane overtake the first plane?
Answer by mananth(16946)

Plane A 280 mph
Plane B = 320 mph
..
difference between their speeds = 320-280 = 40 mph
..
plane A is 6 minutes ahead .
So in 6 minutes it travels 280* 1/10 = 28 miles..
...
Plane B catches up 40 mph per hour
28 miles it will catch up in 28/40 = 7/10 hours
..
plane B speed = 320mph
time to catch up = 7/10 hours
distance = speed * time = 320*7/10
=224 miles
...
m.ananth@hotmail.ca